what are the dimensions of a room that is 1.5 times as long as it is wide, and its perimeter is 25 meters
Found 2 solutions by HyperBrain, rfer:
Answer by HyperBrain(694)

P=2(L + W)
--
Given: L=1.5 W ; P=25 m
---
25=2(1.5 W + W)=2(2.5 W)=5 W
W=25/5=5 m.
L=1.5 (5)=7.5 m.
--
Therefore, the dimensions of the room are 7.5 x 5 m.

Answer by rfer(10432)

2x+2(1.5x)=25
2x+3x=25
x=5
width=5 m
length=7.5 m